# Fault-tolerant hyperbolic Floquet quantum error correcting codes

### v3

@@ Abstract
A central goal in quantum error correction is to reduce the overhead of fault-tolerant quantum computing by increasing noise thresholds and reducing the number of physical qubits required to sustain a logical qubit. We introduce a potential path towards this goal based on a family of dynamically generated quantum error correcting codes that we call “hyperbolic Floquet codes.” These codes are defined by a specific sequence of non-commuting two-body measurements arranged periodically in time that stabilize a topological code on a hyperbolic manifold with negative curvature. We focus on a family of lattices for $n$ qubits that, according to our prescription that defines the code, provably achieve a finite encoding rate $(1/8+2/n)$ and have a depth-3 syndrome extraction circuit. Similar to hyperbolic surface codes, the distance of the code at each time-step scales at most logarithmically in $n$. The family of lattices we choose indicates that this scaling is achievable in practice. We develop and benchmark an efficient matching-based decoder that provides evidence of a threshold near $0.1\,\%$ in a phenomenological noise model and $0.25\,\%$ in an entangling measurements noise model. Utilizing weight-two check operators and a qubit connectivity of 3, one of our hyperbolic Floquet codes uses 400 physical qubits to encode 52 logical qubits with a code distance of 8, i.e., it is a $[[400,52,8]]$ code. At small error rates, comparable logical error suppression to this code requires 5x as many physical qubits (1924) when using the honeycomb Floquet code with the same noise model and decoder.
@@ 1 Introduction
Current schemes either support weight-two check operators or maintain a constant ratio of logical to physical qubits, but not both. For instance, recent research on quantum Low-Density Parity-Check (LDPC) codes [12] suggests that these codes can encode $12$ logical qubits using $288$ physical qubits. This compares favourably against surface codes which would require $4000$ physical qubits to encode $12$ logical qubits for similar error suppression. In an LDPC-type quantum error-correcting code, each check operator acts on only a few qubits, and each qubit is part of only a few checks [20, 21, 22, 23, 24, 12, 25]. However, these particular quantum LDPC codes [12] necessitate the use of weight-six checks, rendering their implementation challenging with current devices. Hyperbolic quantum codes [26, 27, 28, 29] that employ tessellations of closed hyperbolic surfaces have been proposed as a solution for achieving finite encoding rate with weight-three check operators. Recently, there has been considerable attention directed toward a new class of codes known as Floquet codes [13, 30, 31, 32, 33, 34, 35, 36, 37]. These codes feature weight-two check operators but come with the limitation of a vanishing encoding rate; thus, raising the question of whether it is possible to achieve a finite encoding rate using weight-two check operators instead.
@@ 2 The Hyperbolic Floquet Code
Figure 1: A hyperbolic Floquet code with 64 physical qubits. The code is based on an octagonal ({8, 3}) lattice, which is displayed using polar coordinates of its vertices. The 64 physical qubits are located at the vertices of the lattice. Identification of the edges at the boundary makes the lattice a periodic tessellation of a genus-5 manifold. Four such periodic boundary condition identifications are marked by letters $a$-$d$. The octagonal faces are three-colored such that the adjacent faces have different colors. Each edge inherits the color from the two faces it would enter if elongated. Green, blue, and red edges correspond to XX, YY, and ZZ measurement checks, respectively. The blue edges at the ends of the highlighted purple path are identified as a periodic boundary condition, marked with letter $b$. The purple path therefore is a closed path consisting of 10 unique vertices, marked by circles. This closed path hosts a weight-ten noncontractible loop operator, Pauli operators of which are denoted next to its vertices.
The edge colors of the octagonal lattice in Fig. 1 are derived from the face colors, such that each edge inherits its color from the two octagonal faces that it connects and would extend into if elongated. In the Floquet code, these three-colored edges constitute the repetitive rounds of two-body check measurements of the physical qubits of the edges. At rounds ${r=3n}$ (${r=3n+1}$, ${r=3n+2}$) with nonnegative integers $n$, all green (blue, red) checks are measured, as illustrated in Fig. 2(a). Specifically, green, blue, and red checks correspond to measurements of XX, YY, and ZZ Pauli operators on the pairs of physical qubits at the ends of the respective edges. Although any two such measurement checks do not necessarily commute, all the checks within a single round do commute and can be considered to be measured simultaneously.
@@
In this section, we assume errorless evolution of the quantum state of the physical qubits. We further consider that this state starts as a maximally mixed state prior to the round $r=0$ of measurements. Then, after each measurement round, we examine the group of all Pauli operators that have an eigenvalue $+1$ on the current state of the physical qubits, i.e., stabilize the current state. This group is termed the instantaneous stabilizer group (ISG), and before any measurements, consists of solely the identity operator. The check measurements, however, change the structure of the ISG.
@@
After rounds ${r=3n}$ (${r=3n+1}$, ${r=3n+2}$), each XX check (YY check, ZZ check) becomes a generator of the ISG, up to a sign determined by the respective measurement outcome (i.e., if the measurement result of a check is ${-1}$, the negative of that check joins the ISG). For every one of these two-body generators of the ISG, two checks in the subsequent measurement round anticommute with it. As a result, these two-body instantaneous stabilizer generators no longer stabilize the state following the next round. For example, consider the four green (i.e., XX) edges around the red face in Fig. 2(b). After round ${r=3n}$, the two-body checks of these edges, up to their measurement results’ signs, are members of the ISG. The upcoming measurement round, ${r=3n+1}$, is blue (i.e., YY) checks. Each green check anticommutes with two blue checks from of the red face, and hence, will be removed from the ISG upon measurement of any of those blue checks. After round ${r=3n+1}$, each of these blue checks, up to its measurement outcome sign, will be in the ISG.
@@
Although the two-body checks are removed from the ISG following the next measurement round, certain products of them commute with the upcoming measurement checks, and therefore, stay in the ISG. In the example of the red face and green checks in Fig. 2(b), the product of all four green checks commutes with any blue check, ensuring its membership in the ISG after round ${r=3n+1}$. The product of these four green checks can be multiplied with the product of the four blue checks of that red face, deriving a weight-eight plaquette operator. This operator, after round ${r=3n+1}$, is part of the ISG. Since it is equal to a product of X and Y Paulis for each qubit, it is equivalent to a product of Z Paulis (up to a phase). In a similar vein, plaquette operators for green and blue faces are products of X and Y Paulis, respectively. The plaquette operators are depicted in Fig. 2(b).
@@
Figure 2: Measurement protocol and plaquette operators of hyperbolic Floquet codes. (a) One cycle of the Floquet code’s measurement protocol. Each cycle consists of three measurement rounds, each round carried out on XX, YY, and ZZ checks, corresponding to green, blue, and red edges in Fig. 1, respectively. In each measurement, the weight-two Pauli operator of the two qubits at the ends of that edge are measured. (b) Plaquette operators that are products of checks along the edges of each octagon, for green, blue, and red plaquettes. They commute with all check operators on the lattice, and their value can be inferred using the measurement outcomes of the check operators.
@@
Each plaquette operator [Fig. 2(b)] shares the same Pauli operator as the two-body checks intersecting it at a single vertex. Conversely, it has a different Pauli operator than the two-body checks intersecting it at two vertices. Consequently, these plaquette operators commute with all check operators. When measured (during rounds $r=2$, $r=3$, and $r=4$ corresponding to red, green, and blue faces), they will join the ISG, up to a sign derived from the measurements. Beyond round $r=4$, the ISG mod phases reaches a periodic steady state with the logical qubits in a mixed state (i.e., stable against the non-commuting measurements), consistently generated by all plaquette operators and the most recent measurement checks, each acquiring a sign from the measurement outcomes.
@@
Logical operators are the operators that commute with the ISG but are not part of it. Commuting with the ISG is required for the logical information to not be destroyed by the measurement rounds. However, operators that commute with the ISG but are always a part of it, i.e., the plaquette operators and their products, have determined values and cannot be used for restoring information about the logical qubits. The logical operators of the hyperbolic Floquet code are contained within noncontractible closed paths of the hyperbolic lattice. These are the paths that, utilizing the periodic boundary conditions, wrap around the lattice. One closed path of the hyperbolic lattice in Fig. 1 is highlighted in purple.
@@
Just as the product of two-body checks around an octagon commutes with all two-body checks, the product along any loop does the same. We will call these loop operators. The contractible loop operators can be expressed as the product of the plaquette operators within them, and they are part of the ISG beyond round $r=4$. On the other hand, the noncontractible loop operators will never integrate into the ISG. The closed path of one such loop is highlighted with purple in Fig. 1. The displayed periodic boundary condition identifications (edges labeled with letters $a$ to $d$) serve as a reference to confirm that the purple path is indeed closed. The lattice’s periodic boundary conditions identify the two blue edges at the ends of the portrayed path, making the two vertices at each end non-unique. The distinct vertices of the weight-ten loop operator are marked with a circle, and the adjacent Pauli operators represent the components of the loop operator. It is noteworthy that the Pauli operator of each physical qubit is determined by its outgoing edge that is not part of the loop; specifically, X, Y, or Z Paulis for green, blue, or red outgoing edges, respectively. Noncontractible loop operators comprise the Floquet code’s logical operators. As we will discuss below, the number of noncontractible loop operators is twice the genus of the manifold.
@@
Each noncontractible loop operator can be decomposed into the product of two operators, where both individually commute with the elements of the ISG. This decomposition, akin to the ISG, depends on the most recent measurement round. Assume that measurements of the green check operators (XX measurements) have just been carried out (round $r=3n$). The next measurement round is the blue checks (YY measurements). The two operators that commute with the ISG elements and make a noncontractible loop operator are the product of X Paulis on the vertices of the blue edges along its closed path (bottom right panel in Fig. 3) and the product of Y Paulis on the vertices of the green edges along it (top left panel in Fig. 3). These two constructions are generalizable to any measurement round as the most recent measurement Paulis on the edges of the upcoming round and the upcoming measurement Paulis on the edges of the most recent round, and hereinafter will be referred to as the type one and type two logical operators.
@@
Given that, after XX measurements, the type one logical operator has only X Paulis and the type two has two Paulis on each XX check, they commute with the two-body XX checks in the ISG. They also commute with plaquette operators within the ISG. Taking the type one operator as an example, each green and red plaquette shares an integer number of blue edges with the closed path. This means it has an even number of non-identity Paulis in common with the first operator. Moreover, since the path is closed and always leaves any octagon that it enters, an even number of its blue edges are incident at any given blue octagon. Therefore, there are again an even number of mutual non-identity Paulis between the operator and any blue plaquette operator. Similar analysis can be carried out for the type two operator, establishing that both operators commute with the ISG. Further, these operators are independent operators and one of them cannot be obtained by multiplying the other one with some elements of the ISG. These are the two logical operators defined on the noncontractible loop.
@@
So far, the constructed logical operators have been found to commute with the ISG after XX measurements. Notably, this construction ensures that they also commute with the upcoming measurements, specifically YY checks on blue edges. To elaborate, the type one logical operator features two X Paulis on any blue edge along its path and the type two operator is made of Y Paulis. Therefore, YY checks can be measured without destroying these logical operators. However, after YY measurements, some of the upcoming ZZ checks anticommute with the current construction of these logical operators and their measurement is going to affect the logical operators. Nonetheless, before ZZ measurements, we can update the expression of the logical operators by multiplying them with the measured YY checks along their noncontractible loops. The measurement outcomes of these YY checks must be factored into the sign of the updated logical operators.
@@
Consider the type two logical operator (top left panel of Fig. 3). When we multiply this operator by YY checks on the closed path, the Y Paulis shift to the red edges (top middle panel of Fig. 3). Note that this is a type one logical operator after YY (blue) and prior to ZZ (red) measurements. Following the next measurement round (ZZ measurements), we multiply this operator with the measured ZZ checks along the path, making it comprised of X Paulis on the red edges (top right panel of Fig. 3). This is in turn a type two logical operator after ZZ and prior to XX measurements. As seen in Fig. 3, the logical operator switches its type with each measurement round, and reverts to its original expression after six rounds.
@@
Now we focus on finding the lattices that give rise to the particular codes and calculate their code parameters. Each code is based on a periodic octagonal lattice. To identify lattices, we turn to the database of symmetric connected graphs of degree three up to 10,000 vertices, provided in [39]. Within this database, the graphs with a girth of eight qualify as periodic octagonal lattices. These lattices are symmetric (arc-transitive) graphs: for any pair of edges, there exists an automorphism of the graph that maps one edge to the other.
The number of vertices of each lattice determines the number of the physical qubits, $n$, for the corresponding code. An octagonal lattice has a Bravais unit cell comprising 16 vertices, which tessellates the surface of a genus-2 manifold [40]. Thus, every set of 16 vertices in a given octagonal lattice correspond to one unit cell and cover an area equal to $-4\pi$ (in units of negative curvature) on the hyperbolic plane. The genus of the manifold covered by the entire lattice is given by:
@@
We assess the periodic octagonal lattices from [39]. Each lattice in this list that is face-three-colorable is the basis for a hyperbolic Floquet code. As detailed above, the corresponding code parameters, ${[[n,k,d]]}$, can be calculated. Among the codes with a given distance, $d$, the one requiring the fewest physical qubits, $n$, is the most efficient. Table 1 lists the code parameters and lattice properties for the most efficient hyperbolic Floquet codes for even distances from 2 to 10  [^1] The colored adjacency matrices of the hyperbolic and honeycomb lattices used for Floquet codes in this paper, as well as Python scripts for evaluating these codes are available online at https://github.com/fahimniya/HyperbolicFloquetQECC.. The maximum code distance achieved based on octagonal lattices in [39] is 10. We designate these distance-maximizing codes using the prefix ‘H’ followed by their respective number of physical qubits.

### v4

@@ Abstract
A central goal in quantum error correction is to reduce the overhead of fault-tolerant quantum computing by increasing noise thresholds and reducing the number of physical qubits required to sustain a logical qubit. We introduce a potential path towards this goal based on a family of dynamically generated quantum error correcting codes that we call “hyperbolic Floquet codes.” These codes are defined by a specific sequence of non-commuting two-body measurements arranged periodically in time that stabilize a topological code on a hyperbolic manifold with negative curvature. We focus on a family of lattices for $n$ qubits that, according to our prescription that defines the code, provably achieve a finite encoding rate $(1/8+2/n)$ while still requiring only two-body measurements. Similar to hyperbolic surface codes, the distance of the code at each time-step scales at most logarithmically in $n$. The family of lattices we choose indicates that this scaling is achievable in practice. We develop and benchmark an efficient matching-based decoder that provides evidence of a threshold near $0.1\,\%$ in a phenomenological noise model and $0.25\,\%$ in an entangling measurements noise model. Utilizing weight-two check operators and a qubit connectivity of 3, one of our hyperbolic Floquet codes uses 400 physical qubits to encode 52 logical qubits with a code distance of 8, i.e., it is a $[[400,52,8]]$ code. At small error rates, comparable logical error suppression to this code requires 5x as many physical qubits (1924) when using the honeycomb Floquet code with the same noise model and decoder.
@@ 1 Introduction
Current schemes either support weight-two check operators or maintain a constant ratio of logical to physical qubits, but not both. For instance, recent research on quantum Low-Density Parity-Check (LDPC) codes [12] suggests that these codes can encode $12$ logical qubits using $288$ physical qubits. This compares favorably against surface codes which would require $4000$ physical qubits to encode $12$ logical qubits for similar error suppression. In an LDPC-type quantum error-correcting code, each check operator acts on only a few qubits, and each qubit is part of only a few checks [20, 21, 22, 23, 24, 12, 25]. However, these particular quantum LDPC codes [12] necessitate the use of weight-six checks, rendering their implementation challenging with current devices. Hyperbolic quantum codes [26, 27, 28, 29] that employ tessellations of closed hyperbolic surfaces have been proposed as a solution for achieving finite encoding rate with weight-three check operators. Recently, there has been considerable attention directed toward a new class of codes known as Floquet codes [13, 30, 31, 32, 33, 34, 35, 36, 37]. These codes feature weight-two check operators but come with the limitation of a vanishing encoding rate; thus, raising the question of whether it is possible to achieve a finite encoding rate using weight-two check operators instead.
@@ 2 The Hyperbolic Floquet Code
Figure 1: A hyperbolic Floquet code with 64 physical qubits. The code is based on an octagonal ({8, 3}) lattice, which is displayed using polar coordinates of its vertices. The 64 physical qubits are located at the vertices of the lattice. Identification of the edges at the boundary makes the lattice a periodic tessellation of a genus-5 manifold. Four such periodic boundary condition identifications are marked by letters $a$-$d$. The octagonal faces are three-colored such that the adjacent faces have different colors. Each edge inherits the color from the two faces it would enter if elongated. Red, green, and blue edges correspond to XX, YY, and ZZ measurement checks, respectively. The green edges at the ends of the highlighted purple path are identified as a periodic boundary condition, marked with letter $b$. The purple path therefore is a closed path consisting of 10 unique vertices, marked by circles. This closed path hosts a weight-ten noncontractible loop operator, Pauli operators of which are denoted next to its vertices.
The edge colors of the octagonal lattice in Fig. 1 are derived from the face colors, such that each edge inherits its color from the two octagonal faces that it connects and would extend into if elongated. In the Floquet code, these three-colored edges constitute the repetitive rounds of two-body check measurements of the physical qubits of the edges. At rounds r=3n{r=3n} (r=3n+1{r=3n+1}, r=3n+2{r=3n+2}) with nonnegative integers $n$, all red (green, blue) checks are measured, as illustrated in Fig. 2(a). Specifically, red, green, and blue checks correspond to measurements of XX, YY, and ZZ Pauli operators on the pairs of physical qubits at the ends of the respective edges. Although any two such measurement checks do not necessarily commute, all the checks within a single round do commute and can be considered to be measured simultaneously.
@@
In this section, we assume errorless evolution of the quantum state of the physical qubits. We further consider that this state starts as a maximally mixed state prior to the round $r=0$ of measurements. Then, after each measurement round, we examine the group of all Pauli operators that have an eigenvalue +1+1 on the current state of the physical qubits, i.e., stabilize the current state. This group is termed the instantaneous stabilizer group (ISG), and before any measurements, consists of solely the identity operator. The check measurements, however, change the structure of the ISG. Figure 2(c) shows the evolution of the ISG through measurement rounds.
@@
After rounds r=3n{r=3n} (r=3n+1{r=3n+1}, r=3n+2{r=3n+2}), each XX check (YY check, ZZ check) becomes a generator of the ISG, up to a sign determined by the respective measurement outcome (i.e., if the measurement result of a check is 1{-1}, the negative of that check joins the ISG). For every one of these two-body generators of the ISG, two checks in the subsequent measurement round anticommute with it. As a result, these two-body instantaneous stabilizer generators no longer stabilize the state following the next round. For example, consider the four red (i.e., XX) edges around the blue face in Fig. 2(b). After round r=3n{r=3n}, the two-body checks of these edges, up to their measurement results’ signs, are members of the ISG. The upcoming measurement round, r=3n+1{r=3n+1}, is green (i.e., YY) checks. Each red check anticommutes with two green checks from of the blue face, and hence, will be removed from the ISG upon measurement of any of those green checks. After round r=3n+1{r=3n+1}, each of these green checks, up to its measurement outcome sign, will be in the ISG.
@@
Although the two-body checks are removed from the ISG following the next measurement round, certain products of them commute with the upcoming measurement checks, and therefore, stay in the ISG. In the example of the blue face and red checks in Fig. 2(b), the product of all four red checks commutes with any green check, ensuring its membership in the ISG after round r=3n+1{r=3n+1}. The product of these four red checks can be multiplied with the product of the four green checks of that blue face, deriving a weight-eight plaquette operator. This operator, after round r=3n+1{r=3n+1}, is part of the ISG. Since it is equal to a product of X and Y Paulis for each qubit, it is equivalent to a product of Z Paulis (up to a phase). In a similar vein, plaquette operators for red and green faces are products of X and Y Paulis, respectively. The plaquette operators are depicted in Fig. 2(b).
@@
Figure 2: Measurement protocol and plaquette operators of hyperbolic Floquet codes. (a) One cycle of the Floquet code’s measurement protocol. Each cycle consists of three measurement rounds, each round carried out on XX, YY, and ZZ checks, corresponding to red, green, and blue edges in Fig. 1, respectively. In each measurement, the weight-two Pauli operator of the two qubits at the ends of that edge are measured. (b) Plaquette operators that are products of checks along the edges of each octagon, for red, green, and blue plaquettes. They commute with all check operators on the lattice, and their value can be inferred using the measurement outcomes of the check operators. (c) The instantaneous stabilizer group (ISG) after each measurement round expressed as generated by its generators after each measurement round. The generators can be check operators (panel a) or plaquette operators (panel b). Each check or plaquette operator represents all similar operators with the same color up to a sign.
@@
Each plaquette operator [Fig. 2(b)] shares the same Pauli operator as the two-body checks intersecting it at a single vertex. Conversely, it has a different Pauli operator than the two-body checks intersecting it at two vertices. Consequently, these plaquette operators commute with all check operators. When measured (during rounds $r=2$, $r=3$, and $r=4$ corresponding to blue, red, and green faces), they will join the ISG, up to a sign derived from the measurements. Beyond round $r=4$, the ISG mod phases reaches a periodic steady state with the logical qubits in a mixed state (i.e., stable against the non-commuting measurements), consistently generated by all plaquette operators and the most recent measurement checks, each acquiring a sign from the measurement outcomes.
@@
Logical operators are the operators that commute with the ISG but are not part of it. Commuting with any two consecutive instances of the ISG is required for the logical information to not be destroyed by the measurement rounds [39]. However, operators that commute with the ISG but are always a part of it, i.e., the plaquette operators and their products, have determined values and cannot be used for restoring information about the logical qubits. The logical operators of the hyperbolic Floquet code are contained within noncontractible closed paths of the hyperbolic lattice. These are the paths that, utilizing the periodic boundary conditions, wrap around the lattice. One closed path of the hyperbolic lattice in Fig. 1 is highlighted in purple.
@@
Just as the product of two-body checks around an octagon commutes with all two-body checks, the product along any loop does the same. We will call these loop operators. The contractible loop operators can be expressed as the product of the plaquette operators within them, and they are part of the ISG beyond round $r=4$. On the other hand, the noncontractible loop operators will never integrate into the ISG. The closed path of one such loop is highlighted with purple in Fig. 1. The displayed periodic boundary condition identifications (edges labeled with letters $a$ to $d$) serve as a reference to confirm that the purple path is indeed closed. The lattice’s periodic boundary conditions identify the two green edges at the ends of the portrayed path, making the two vertices at each end non-unique. The distinct vertices of the weight-ten loop operator are marked with a circle, and the adjacent Pauli operators represent the components of the loop operator. It is noteworthy that the Pauli operator of each physical qubit is determined by its outgoing edge that is not part of the loop; specifically, X, Y, or Z Paulis for red, green, or blue outgoing edges, respectively. Noncontractible loop operators comprise the Floquet code’s logical operators. As we will discuss below, the number of noncontractible loop operators is twice the genus of the manifold.
@@
Each noncontractible loop operator can be decomposed into the product of two operators, where both individually commute with the elements of the ISG. This decomposition, akin to the ISG, depends on the most recent measurement round. Assume that measurements of the red check operators (XX measurements) have just been carried out (round $r=3n$). The next measurement round is the green checks (YY measurements). The two operators that commute with the ISG elements and make a noncontractible loop operator are the product of X Paulis on the vertices of the green edges along its closed path (bottom right panel in Fig. 3) and the product of Y Paulis on the vertices of the red edges along it (top left panel in Fig. 3). These two constructions are generalizable to any measurement round as the most recent measurement Paulis on the edges of the upcoming round and the upcoming measurement Paulis on the edges of the most recent round, and hereinafter will be referred to as the type one and type two logical operators.
@@
Given that, after XX measurements, the type one logical operator has only X Paulis and the type two has two Paulis on each XX check, they commute with the two-body XX checks in the ISG. They also commute with plaquette operators within the ISG. Taking the type one operator as an example, each red and blue plaquette shares an integer number of green edges with the closed path. This means it has an even number of non-identity Paulis in common with the first operator. Moreover, since the path is closed and always leaves any octagon that it enters, an even number of its green edges are incident at any given green octagon. Therefore, there are again an even number of mutual non-identity Paulis between the operator and any green plaquette operator. Similar analysis can be carried out for the type two operator, establishing that both operators commute with the ISG. Further, these operators are independent operators and one of them cannot be obtained by multiplying the other one with some elements of the ISG. These are the two logical operators defined on the noncontractible loop.
@@
So far, the constructed logical operators have been found to commute with the ISG after XX measurements. Notably, this construction ensures that they also commute with the upcoming measurements, specifically YY checks on green edges. To elaborate, the type one logical operator features two X Paulis on any green edge along its path and the type two operator is made of Y Paulis. Therefore, YY checks can be measured without destroying these logical operators. However, after YY measurements, some of the upcoming ZZ checks anticommute with the current construction of these logical operators and their measurement is going to affect the logical operators. Nonetheless, before ZZ measurements, we can update the expression of the logical operators by multiplying them with the measured YY checks along their noncontractible loops. The measurement outcomes of these YY checks must be factored into the sign of the updated logical operators.
@@
Consider the type two logical operator (top left panel of Fig. 3). When we multiply this operator by YY checks on the closed path, the Y Paulis shift to the blue edges (top middle panel of Fig. 3). Note that this is a type one logical operator after YY (green) and prior to ZZ (blue) measurements. Following the next measurement round (ZZ measurements), we multiply this operator with the measured ZZ checks along the path, making it comprised of X Paulis on the blueedges (top right panel of Fig. 3). This is in turn a type two logical operator after ZZ and prior to XX measurements. As seen in Fig. 3, the logical operator switches its type with each measurement round, and reverts to its original expression after six rounds.
@@
Now we focus on finding the lattices that give rise to the particular codes and calculate their code parameters. Each code is based on a periodic octagonal lattice. To identify lattices, we turn to the database of symmetric connected graphs of degree three up to 10,000 vertices, provided in [40] [^1] Reference [40] includes methodology and graphs up to 768 vertices. Graphs with up to 10,000 vertices can be found here.. Within this database, the graphs with a girth of eight qualify as periodic octagonal lattices. These lattices are symmetric (arc-transitive) graphs: for any pair of edges, there exists an automorphism of the graph that maps one edge to the other.
The number of vertices of each lattice determines the number of the physical qubits, $n$, for the corresponding code. The number of logical qubits, $k$, on the other hand, is determined by the genus of the manifold that the periodic lattice tessellates. In order to find the relation between the genus of this manifold and $n$, we need to look at the crystallography structure of the octagonal lattice [41].
An octagonal lattice has a Bravais unit cell comprising 16 vertices, which tessellates the surface of a genus-2 manifold (see Fig. 12 in Ref. [41]). Every additional set of 16 vertices in a given octagonal lattice corresponds to one unit cell and cover an area equal to $4\pi$ (in units of negative curvature) on the hyperbolic plane. Thus, the genus of the manifold covered by the entire lattice is given by:
@@
We assess the periodic octagonal lattices from [40]. Each lattice in this list that is face-three-colorable is the basis for a hyperbolic Floquet code. As detailed above, the corresponding code parameters, [[n,k,d]]{[[n,k,d]]}, can be calculated. Among the codes with a given distance, $d$, the one requiring the fewest physical qubits, $n$, is the most efficient, i.e., it uses the least amount of resources to achieve the same distance. Table 1 lists the code parameters and lattice properties for the most efficient hyperbolic Floquet codes for even distances from 2 to 10  [^2] The colored adjacency matrices of the hyperbolic and honeycomb lattices used for Floquet codes in this paper, as well as Python scripts for evaluating these codes are available online at https://github.com/fahimniya/HyperbolicFloquetQECC.. The maximum code distance achieved based on octagonal lattices in [40] is 10. We designate these distance-maximizing codes using the prefix ‘H’ followed by their respective number of physical qubits.